OpenStudy (anonymous):

A triangle has sides with lengths of 6 yards, 13 yards, and 17 yards. Is it a right triangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you give me a medal?

OpenStudy (fortytherapper):

How is it a right triangle? I'll vote "No"

OpenStudy (sleepyhead314):

Hello @mia1902 and Welcome to OpenStudy! :) Sorry for the late reply, but to clear this up... Right triangles all follow the Pythagorean Theorem where \(a^2 + b^2 = c^2\) in other words, one of the legs squared plus the other leg squared is equal to the square of the longest side. In this case, \(if\) this triangle was a right triangle, then \(6^2 + 13^2 = 17^2\) should be true. \(if\) you end up with something weird (not true), then the triangle would Not be a right triangle I hope I was able to help ^_^ If you need me to explain more, feel free to tag me with @sleepyhead314 :)